cwf0:

(05.03 HC) There are 7 tigers at the zoo. Each year, the number of tigers increases by 10%. There are 15 eagles at the zoo. Each year, 2 new eagles join the zoo. Part A: Write functions to represent the number of tigers and eagles at the zoo throughout the years. (4 points) Part B: How many tigers are at the zoo after 10 years? How many eagles are at the zoo after the same number of years? (2 points) Part C: After approximately how many years is the number of tigers and eagles the same? Justify your answer mathematically. (4 points) (10 points)

jhonyy9:

what mean this 10% in numbers in case of tigers and what percent mean this 2 in case of eagles ?

jhonyy9:

ok so to calcule the 10% of 7 tigers use the rule of 3 simple and in this way you write 7 tigers mean 100% x tigers mean 10% cross multiplie and calcule the value of x x = (7*10)/100 = 70/100 = 0,7 so from this result that the number of tigers in this zoo increase by 7 tigers in 10 years do you can continue with eagles ?